1.) Solve by the elimination method:
7r-9s=-35
9r+7s=85
2.) Solve for the following system of equations:
x+4y=2
x=8-4y
3.)Find the slope and the y-intercept: f(x)=-2x-3
4.)Solve. 0.9x+9<_1.7x-8

Answer by richwmiller(17219)

You are allowed only one problem at a time
#2
x+4y=2
x=8-4y
rearrange
x+4y=2
x+4y=8
they are parallel so there is no solution